OpenStudy (anonymous):

find the derivative of x^3 * cosX ...would like to compare my steps with yours please

OpenStudy (solomonzelman):

use the product rule./

OpenStudy (anonymous):

used product rule first

OpenStudy (solomonzelman):

okay, and what were you getting?

OpenStudy (anonymous):

3x^2(cosx) + x^3(-sinX)

OpenStudy (solomonzelman):

yes that is correct. you can re-write it a bit better though.

OpenStudy (anonymous):

factor x^2?

OpenStudy (solomonzelman):

well, \(\large\color{black}{ 3x^2(\cos x) + x^3(-\sin x) }\) is same as, \(\large\color{black}{ 3x^2 \cos x - x^3 \sin x }\)

OpenStudy (anonymous):

is XsinX an identity?

OpenStudy (solomonzelman):

you don't need to factor. If that is what your teacher would required or (even) prefer, then you can do that, that is not a problem:)

OpenStudy (solomonzelman):

I don't think there is really an identity for \(\large\color{black}{ x\sin x }\).

OpenStudy (anonymous):

x^2(3cosX - XsinX)

OpenStudy (solomonzelman):

yes:) you can do that too, I don't see a need though, unless you are going to be finding the critical numbers.

OpenStudy (anonymous):

what would the equation for the tangent line at x=pi be?

OpenStudy (solomonzelman):

you will need to plug in pi, into the derivative. this will give you the slope. then the point is (pi, f(pi) )

OpenStudy (solomonzelman):

\(\large\color{black}{ f(x)=x^3\cos x}\) \(\large\color{black}{ f(\pi)=(\pi)^3\cos (\pi)}\) \(\large\color{black}{ f(\pi)=(\pi)^3(-1)}\) \(\large\color{black}{ f(\pi)=-(\pi)^3}\)

OpenStudy (solomonzelman):

so your point is, \(\large\color{black}{ (\pi,~-(\pi)^3~)}\)

OpenStudy (solomonzelman):

then \(\large\color{black}{ f \prime(\pi)}\) is the slope.

OpenStudy (solomonzelman):

We had: \(\large\color{black}{ f\prime(x)=x^2(3 \cos x - x \sin x)}\)

OpenStudy (anonymous):

for the answer i get

OpenStudy (solomonzelman):

can you find, \(\large\color{black}{ f\prime(\pi)}\) ?

OpenStudy (anonymous):

the equation at pi is 3(pi)^2x - 4(pi)^3

OpenStudy (solomonzelman):

I found the point for you, and you need to find the slope. then apply the point slope formula.

OpenStudy (anonymous):

3((pi)^2)(x) - 4(pi)^3

OpenStudy (solomonzelman):

I don't understand what yu aregetting, and after plugging pi for x, why do you still have an x?

OpenStudy (anonymous):

equation of the tangent line at that point

OpenStudy (solomonzelman):

equation has " = "

OpenStudy (solomonzelman):

\(\large\color{black}{ f\prime(\pi)=(\pi)^2(3 \cos\pi- \pi \sin \pi)}\) \(\large\color{black}{ f\prime(\pi)=(\pi)^2(3 \cos\pi- \pi \sin \pi)}\) \(\large\color{black}{ f\prime(\pi)=(\pi)^2(3 \cos\pi)}\) \(\large\color{black}{ f\prime(\pi)=(\pi)^2(3)(-1)}\) \(\large\color{black}{ f\prime(\pi)=-3(\pi)^2}\)

OpenStudy (solomonzelman):

\(\large\color{black}{ f\prime(\pi)=m=-3(\pi)^2}\)

OpenStudy (solomonzelman):

\(\large\color{black}{ (y-y_1)=m(x-x_1)}\)

OpenStudy (anonymous):

actually got -3(pi^2)x + 2(pi^3)

OpenStudy (solomonzelman):

got for what?

OpenStudy (anonymous):

the slope of the tangent line

OpenStudy (anonymous):

y = -3(pi^2)x + 2(pi^3), sorry, the equation of the tangent line

OpenStudy (solomonzelman):

slope: \(\large\color{black}{ f\prime(\pi)=-3(\pi)^2 }\) point: \(\large\color{black}{ (\pi,-\pi^3) }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ y-(-\pi^3)=-3\pi^2~(x-\pi) }\) \(\large\color{black}{ y-(-\pi^3)=-3\pi^2x+3\pi^3 }\) \(\large\color{black}{ y+\pi^3=-3\pi^2x+3\pi^3 }\) \(\large\color{black}{ y=-3\pi^2x+2\pi^3 }\)

OpenStudy (solomonzelman):

you were correct, I was just checking that.

OpenStudy (anonymous):

awesome, for some reason, wolfram alpha spits out a different answer for our derivative, but it looks like we did it correctly, maybe it's their algorithm?

OpenStudy (anonymous):

you think we're right, though?